### v2

# pasta: Pattern Analysis for Spatial Omics Data

Abstract
Spatial omics assays allow for the molecular characterisation of cells in their spatial context. Notably, the two main technological streams, imaging-based and high-throughput sequencing-based, can give rise to very different data modalities. The characteristics of the two data types are well known in adjacent fields such as spatial statistics as point patterns and lattice data, and there is a wide range of tools available. This paper discusses the application of spatial statistics to spatially-resolved omics data and in particular, discusses various advantages, challenges, and nuances. This work is accompanied by a vignette, pasta (https://robinsonlabuzh.github.io/pasta), that showcases the usefulness of spatial statistics in biology using several R and Python packages.
Introduction
Molecular profiling of cells in organs or tissues can be accomplished either in bulk or at the single-cell level. However, tissue must be dissociated, which may select against certain cell types and results in the loss of the spatial organization of the cells. This limitation is overcome with spatial profiling techniques, which can range from spatial proteomics (e.g., IMC, MIBI-TOF, 4i, CODEX, mIF [giesenHighlyMultiplexedImaging2014, kerenMIBITOFMultiplexedImaging2019, gutMultiplexedProteinMaps2018, blackCODEXMultiplexedTissue2021, tsujikawaQuantitativeMultiplexImmunohistochemistry2017]), to spatial transcriptomics assays based on fluorescence in situ hybridisation (e.g., FISH, MERFISH, CosMx, Xenium [langer-saferImmunologicalMethodMapping1982, chenSpatiallyResolvedHighly2015, janesickHighResolutionMapping2023]) and based on sequencing (e.g., 10x Visium and Visium HD, Slide-seq V1&V2 [stahlVisualizationAnalysisGene2016, oliveiraCharacterizationImmuneCell2024, rodriquesSlideseqScalableTechnology2019, stickelsHighlySensitiveSpatial2021]) to even spatial epigenomics [luSpatiallyResolvedEpigenomic2022]. Moreover, the combination of individual methods enabled spatial multiomic research [vandereykenMethodsApplicationsSinglecell2023]. The technologies and their application in biological research have been the topic of various detailed reviews [aspSpatiallyResolvedTranscriptomes2020, moffittEmergingLandscapeSpatial2022, raoExploringTissueArchitecture2021, mosesMuseumSpatialTranscriptomics2022, pallaSpatialComponentsMolecular2022, bressanDawnSpatialOmics2023].
In this work, we focus on the concepts and application of exploratory statistical approaches for spatial data with a focus on downstream steps; our overview also centres around key differences between data modalities that result from current spatial profiling technologies. In this regard, we complement the ongoing discussion [zormpasMappingTranscriptomeRealizing2023, mosesVoyagerExploratorySinglecell2023] about the challenges in applying spatial statistics to spatial (transcript)omics data.
Most spatial omics assays can be classified as either high-throughput sequencing (HTS) based or imaging-based. In HTS-based approaches, positional information is recorded using a predetermined array of spots or beads. Imaging-based approaches, however, either target the molecules of interest with fluorescent probes, ablate regions stained with a cocktail of antibodies via metal tag readouts, or amplify and sequence target sequences in situ. Several technologies are emerging, but the main trade-offs stem from the resolution, number of features, and sensitivity of the readout [raoExploringTissueArchitecture2021, parkSpatialOmicsTechnologies2022]. For example, HTS-based approaches capture the entire transcriptome (i.e., untargeted), but come at a resolution determined by the spot size. On the other hand, imaging-based approaches typically have lower depth of information (i.e., targeted) but show a higher resolution in comparison to HTS-based approaches. Note also that the landscape of spatially-resolved molecular technologies is rapidly changing, mostly in terms of higher resolution for HTS-based methods and comprehensive marker panels for imaging-based approaches (Figure 1, Technology) [moffittEmergingLandscapeSpatial2022, raoExploringTissueArchitecture2021, rademacherComparisonSpatialTranscriptomics2024].
In terms of data analysis, the technological streams are quite distinct. HTS-based approaches collect data along regularly-spaced spots or beads resulting in a so-called regular lattice. In contrast, imaging-based approaches measure features at exact locations that can be assumed to have originated from a stochastic process known as a point process [raoExploringTissueArchitecture2021]. Thus, we will distinguish between lattice-based and point pattern-based spatial omic data. While the focus is often on the technology (c.f., imaging-based versus HTS-based), we argue that the distinction should lie at the data representation level [raoExploringTissueArchitecture2021]. This is important, since there are technologies that can be represented as both point patterns and lattice data, depending on the resolution and data processing. For example, imaging-based technologies that detect transcripts at subcellular resolution result in data where transcript locations can be represented as a point pattern, while for the same dataset, segmented cells can be represented either as an irregular lattice, or as a point pattern via their cell centroids. In general, if events of a point pattern are aggregated in specified regions (e.g., transcripts per cell from segmentation), we end up with lattice data (Figure 1, Data modality) [daleSpatialAnalysisGuide2014, pebesmaSpatialDataScience2023].
Since we can represent some spatial omic datasets as either lattice or point pattern, it is important to further understand the assumptions and analysis strategies of such data modalities [cressieStatisticsSpatialData1993]. On the one hand, spatial point pattern analysis assumes that the point locations were generated by a stochastic process, a so-called event-based view of the data; the goal is to study the properties of this process [baddeleySpatialPointPatterns2015]. On the other hand, one could assume that the locations are fixed and known at the time of sampling and study the associated features at each location via an observation-based view of the data, while recognising that the observations in the lattice are not independent due to the spatial structure [anselinWhatSpecialSpatial1989].
Spatial data modalities have been present in other fields for decades, such as geography, and in particular, the field of spatial statistics offers a large analysis toolbox. For example, spatial omics data collected across a lattice can be analysed by exploratory tools, such as Local Moran’s and the bivariate Lee’s statistic. Imaging data generated by a stochastic process can be input for methods of point pattern analysis, including summaries such as Ripley’s or empty-space functions [zuurAnalysingEcologicalData2007, daleSpatialAnalysisGuide2014, baddeleySpatialPointPatterns2015, pebesmaSpatialDataScience2023] (Figure 1, Analysis).
There are already several tools for omics analysis that employ spatial statistics approaches. Amongst others, Voyager is a framework and a collection of use cases for lattice data, written in both R and Python [mosesVoyagerExploratorySinglecell2023]; it provides data structures and functionalities to compute spatial statistics in molecular biology. Voyager comes with comprehensive vignettes for several spatial profiling technologies. The methodological focus of their vignettes is on lattice data analysis and the spdep geospatial package. Since Voyager offers efficient implementations for lattice data analysis in R, we build on their framework in our vignettes for the lattice data component. Other methods written in R include: SPIAT, which contains various tools for spatial analysis, such as: neighbourhood analysis, local metrics and heterogeneity scores etc. [fengSpatialAnalysisSPIAT2023]; scFeatures, a toolbox comprising lattice data analysis functionality such as Moran’s and point pattern metrics as the -function [caoScFeaturesMultiviewRepresentations2022]; spicyR, which uses a compressed version of the global -function for cross-sample comparison [caneteSpicyRSpatialAnalysis2022]; lisaClust, a spatial domain detection method via LISA (local indicators of spatial association) curves [patrickSpatialAnalysisHighly2023]; mxfda, that uses point pattern summary functions for survival analysis [wrobelMxfdaComprehensiveToolkit2024]; spatialDM uses global and local bivariate Moran’s to score colocalisation of ligand-receptor pairs [liSpatialDMRapidIdentification2023]; and MERINGUE, which uses nearest-neighbour autocorrelation for spatially variable gene selection [millerCharacterizingSpatialGene2021]. Methods in Python include squidpy, a package with various spatial statistics tools for both point pattern and lattice data [pallaSquidpyScalableFramework2022] and PySAL a general Python library for spatial analysis [reyPySALPythonLibrary2010]. For a detailed view on general methods development in the field of spatial transcriptomics, we refer readers to the online Supplementary table of the museum of spatial transcriptomics paper [mosesMuseumSpatialTranscriptomics2022].
Here, we will explore the application of spatial statistics to spatially-resolved omics data guided by the two main streams of data, point pattern-based and lattice-based methods. We will compare the two streams and show their strengths and limitations. Finally, we give an outlook on the challenges and research gaps in the field of spatial data analysis for omics data. Furthermore, we provide a vignette (https://robinsonlabuzh.github.io/pasta) that showcases analysis of data from multiple technologies, where concepts and assumptions are discussed in detail, with inline R and Python code. Overall, our resource Pattern Analysis for SpaTial omics datA (pasta) will highlight the usefulness and transferability of existing exploratory spatial statistics approaches in the context of spatial tissue profiling.
Results
Point Pattern Analysis
Definitions
Most imaging-based technologies give a high-resolution readout of either subcellular compartments (e.g., laser ablation pixels in IMC [giesenHighlyMultiplexedImaging2014]) or even individual transcripts (e.g., FISH [feminoVisualizationSingleRNA1998]). In molecule-level technologies, the expression of features (transcript quantifications, ion counts, fluorescent intensities) is recorded where they occur [aspSpatiallyResolvedTranscriptomes2020]. Often, we are interested in the distribution and composition of cell types. In order to annotate cells, we require (good) cell segmentation boundaries (an active area of research not discussed here; see [wangCellSegmentationImage2019, stringerCellposeGeneralistAlgorithm2021, greenwaldWholecellSegmentationTissue2022, petukhovCellSegmentationImagingbased2022, architSegmentAnythingMicroscopy2023]), allowing transcripts to be allocated to a given cell [fuBIDCellBiologicallyinformedSelfsupervised2024], and then cell type inference. Since the locations of the annotated cells are generated by a stochastic biological process, we can approximate e.g., the cell centroids as points and analyse their location as point patterns.
In the analysis of these point patterns, the goal is to make inferences on the point process that generated the data, not on the patterns themselves [baddeleySpatialPointPatterns2015, p. 127]. Points in a point pattern can also carry a mark, which is a (univariate or multivariate) variable associated with a point [baddeleySpatialPointPatterns2015, pp. 147, 563]. In cell biology, we could imagine both categorical (e.g., cell type or spatial domain labels) or continuous marks (e.g., gene expression).
When a point pattern with many cell types is considered, there are two approaches to formulate the setup. In the first variant, we consider cells in a tissue to depend on each other so their distribution is due to one overarching biological process; this is referred to as a multitype view. In contrast, when we consider the patterns to be created by point processes, we assume that these processes can be individually analysed; this represents the multivariable viewpoint (change in terminology in comparison to [baddeleySpatialPointPatterns2015]). The processes do not need to be independent, but the independence of components (and functions) is often used as a null model for multivariable analysis (see box Assumptions) [baddeleySpatialPointPatterns2015, pp. 565ff.].
Assumptions
Many imaging platforms measure patterns in a field of view (FOV) and collect several such FOVs per sample. In some technologies, there are no gaps between the individual FOVs and they can be assembled into one big image. If there are gaps, consecutive FOVs can either be stitched together (which can be computationally complex) or analysed individually [aspSpatiallyResolvedTranscriptomes2020, mohammadiFastRobustFeaturebased2024]. When analysing individual FOVs, we do not observe the entire pattern but rather a subset of the bigger pattern. This is called window sampling because the window is just a sample of the bigger point pattern. A related concept is the small world model, which describes that points can only be observed in a finite world and not beyond these boundaries [baddeleySpatialPointPatterns2015, pp. 143ff.]. Often, the distinction between the window sampling and small world model concepts is not clear. An example is the arrangement of the epithelium, which can can be imaged with several FOVs (window sampling) but there are no cells expected outside of the epithelial boundary (small world) [garrido-trigoMacrophageNeutrophilHeterogeneity2023]. Therefore, in cell biology, we often encounter a mixture of these two concepts.
Apart from the window of measurement, we need to make assumptions on the statistical properties of the point pattern, including, most importantly, whether points can be considered homogeneous or not. Homogeneity (see Box Assumptions) assumes that the number of points in a given region is proportional to its area , for arbitrary ; i.e., homogeneity refers to a uniform intensity of points across the window of measurement. If that is not the case, the point process is said to be inhomogeneous [baddeleySpatialPointPatterns2015, p.132-133]. This difference has important implications for the analysis of a process. For example, when adjusting for varying local intensities, the process could be found to be inhomogeneous rather than genuinely clustered. This is called the confounding between intensity and interaction (see box Assumptions) [baddeleySpatialPointPatterns2015, pp.151-152].
Options for analysis
Point pattern analysis offers a broad spectrum of options. Most often, we compare patterns in 2D (e.g., a single histological slice), but the methods are applicable with some adjustments to 3D patterns (e.g., confocal microscopy volumes) [baddeleySpatialPointPatterns2015, p. 651], with some increased computational cost [baddeleyAnalysisThreeDimensionalPoint1993].
As mentioned above, points can carry marks; the nature of these marks determines what options are available to the analyst. For example, many point pattern methods are designed for categorical data, such as cell types or spatial domain labels [baddeleySpatialPointPatterns2015, p.147] and can be adapted to the analysis of continuous marks [baddeleySpatialPointPatterns2015, pp.644 ff.]. Model-based approaches for spatial data might also be of interest. This is an entire subfield on its own with many options available [baddeleySpatialPointPatterns2015, pp.297 ff.][pebesmaSpatialDataScience2023].
When comparing categorical marks, the analyst must decide whether to compare a single type (within-type comparison) or between a pair of types (cross-type comparison). For example, a within-type comparison could use the pattern of epithelial cells in a microscopy image; the correlation, spacing or distances between epithelial cells could be summarised. In contrast, a cross-type analysis could consider the correlation between epithelial cells and B cells [baddeleySpatialPointPatterns2015, pp. 561ff.].
Examples
A well-known point pattern summary is Ripley’s -function, defined as the expected number of points falling in a specified radius around an arbitrary point, averaged over all points in the observation window (e.g., FOV) (potentially adjusted for edge effects; see box Assumptions). Ripley’s -function thus quantifies the correlation structure of a point pattern, with within-type and cross-type versions. A point pattern where points have no preference for location shows complete spatial randomness (CSR). This means that the number of points that fall in an arbitrary region follow a Poisson distribution with a rate parameter that is proportional to the size of the region. A clustered process would have values of the -function larger than the Poisson process at that specific scale; values of smaller than the Poisson process indicate a self-inhibiting process (i.e., where points repel each other) [ripleySecondOrderAnalysisStationary1976, ripleyModellingSpatialPatterns1977][baddeleySpatialPointPatterns2015, pp.203 ff.].
Ripley’s -function can be variance stabilised by taking the square root, known as Besag’s -function [besag1977contribution]. It is a global estimate that measures correlation across an entire observation window and can be used instead of the standard -function [baddeleySpatialPointPatterns2015, p.402], [baddeleySpatstatPackageAnalyzing2005].
Figure 2A-C show point patterns of mature oligodendrocytes over three different serial sections of the same mouse hypothalamic preoptic region (the numbers , , indicate the -axis position of the slices: one slice every 50 m along the anterior-posterior axis [moffittMolecularSpatialFunctional2018]). Figure 2D shows the homogeneous global -function, which quantifies the correlation within each slice and the respective -function for a Poisson process (dashed line). The homogeneous -function indicates strong clustering, even for the slice, however, the inhomogeneous -function, which makes a local adjustment for intensity, indicates a level of homogeneity very similar to a (completely spatially random) Poisson process (Supp. Figure S1A-C vs. D-F). This is an example of intensity vs. interaction confounding (see box Assumptions). If slice should therefore be interpreted as clustered, depends on the assumption of homogeneity for that slice. In addition, there are differences between the inhomogeneity corrections in Figure S3; the corrections are discussed in more detail in the vignette.
Local versions of the various summary functions exist. The LISA (local indicators of spatial association) framework provides local versions for many spatial statistics (including point pattern and lattice data analysis) [anselinLocalIndicatorsSpatial1995]. For example, Figure 2E shows LISA curves for the slice [getisSecondOrderNeighborhoodAnalysis1987, anselinLocalIndicatorsSpatial1995]. That is, for each point, we calculate a local -function. These curves can be analysed with functional principal component analysis (fPCA), which allows to extract the main functional modes of variation [ramsayPrincipalComponentsAnalysis2005]. The idea to perform fPCA on spatial statistics summaries has been used in ecological and spatial omics data [illianPrincipalComponentAnalysis2006, vuFunSpaceFunctionalSpatial2023, vuSPFSpatialFunctional2022]. Figure 2F shows the scores of the first two functional PCs. There are two main clusters, representing either the physically clustered oligodendrocytes or those in low-density regions.
Challenges and limitations
Analysing biological samples formally as point patterns comes with several challenges and limitations. The first and most important is viewing the biological sample as the realisation of a point process. It is generally not easy to define whether the entire pattern of several cell types should be represented as a single point process (multitype analysis). Analysing the cell types individually and considering them as separate point processes might underestimate dependencies among the cells (multivariable analysis) [baddeleySpatialPointPatterns2015, pp. 565 ff.]. If, as in our example in Figure 2, the sections are far apart from each other, it can be justified to analyse the processes in isolation as a multivariable analysis. However, if we consider cells in two adjacent slices, these have probably been formed by one biological process and should not be analysed in isolation.
Obtaining a representative sample of a point pattern is also a challenge. In spatial omics, we are often provided with FOVs that an experimentalist has selected from a much larger region or that stem from the technology itself. Such a sample might have different characteristics compared to the point pattern of the larger region, as FOVs are sometimes selected based on morphological properties (e.g., H&E staining). In cases where we have a small world scenario, we would want to limit our observation window to the full region where the points can occur.
Lattice Data Analysis
Definitions
In contrast to point pattern-based methods, analysts could take the observation-based view and regard the location of cells or spots as a fixed (regular or irregular) lattice and consider the corresponding marker expression at each cell or spot. Several HTS-based spatial transcriptomics technologies produce data on a regular lattice (i.e., evenly spaced spots or beads of uniform size and shape), whereas imaging-based technologies yield irregular lattice structures (i.e., segmentation resulting in variable cell sizes and shapes, and non-uniform spacing) [zuurAnalysingEcologicalData2007, pp. 321ff.] [anselin2024introduction].
Assumptions
Once we have recorded the outline and arrangement of the spatial units in the lattice, we can specify the strength of spatial relationships between each unit. Each pair of units is assigned a weight: the stronger the connection between units, the higher the weight. The collection of all weights between locations and , , forms the weight matrix (see box Options for analysis). The construction of the weight matrix is critical for all downstream analyses as it encodes the spatial relationship between the units in the lattice [zuurAnalysingEcologicalData2007, pp. 321ff.] and there exist various strategies: contiguity-based (i.e., in direct contact), graph-based, distance-based, or higher-order neighbours [getisSpatialWeightsMatrices2009, zuurAnalysingEcologicalData2007, pebesmaSpatialDataScience2023].
Options for analysis
A common exploratory analysis for lattice data is based on the concept of spatial autocorrelation. Spatial autocorrelation measures the degree of association of features, e.g., the expression of a gene, that are assumed to be dependent in space. For each combination of spatial locations, a measure of association is calculated and scaled by the weight of the connection [anselin2024introduction]. This approach considers both the proximity (via the weight matrix) and the characteristics of the locations (via the metric) [zuurAnalysingEcologicalData2007, pp. 327ff.] [pebesmaSpatialDataScience2023, pp. 209ff.] [getisAnalysisSpatialAssociation1992]. Over the years, various association measures have emerged, each specific to an aspect of the data. While many allow for univariate comparisons, in which one variable (e.g., expression of a gene) is compared over multiple locations, others allow for the comparison of two (bivariate) or more (multivariate) features at once [pebesmaSpatialDataScience2023, mosesVoyagerExploratorySinglecell2023]. For most metrics, the features are assumed to be continuous measurements. However, metrics allow for the measurement of spatial autocorrelation of categorical variables (e.g., join count statistics). In addition to the number of features and the type of measurement that can be compared in the association score, there is the to consider: global spatial autocorrelation metrics estimate the average level of spatial autocorrelation across all locations whereas local measures give a statistic at each location. The global statistic can be seen as the weighted sum of its respective local statistics [anselinLocalIndicatorsSpatial1995].
Example
Moran’s is a prominent example of a spatial autocorrelation metric. It can be interpreted as the mean Pearson correlation between values at locations and the mean value of the corresponding neighbours [moranNotesContinuousStochastic1950]. A positive value of Moran’s indicates that locations with similar values are clustered while a negative value indicates the clustering of dissimilar values, giving a measure of spatial heterogeneity. The global value of Moran’s is bounded by and with an expected value under spatial randomness near for large , (). Figure 3A shows evidence for spatial autocorrelation of Nrgn expression in mouse brain tissue; the (global) Moran’s value is (p-value essentially zero). Global measures can be used to identify spatially variable genes when compared to the global autocorrelation values of different genes [weberNnSVGScalableIdentification2023, chenEvaluatingSpatiallyVariable2024]. Furthermore, there is a local version of Moran’s (interpretation is analogous to the global version). Figure 3B shows the estimate of local Moran’s at each measured spot (n.b., the local version is not bounded by and ). Figure 3C shows the corresponding adjusted p-values (testing for non-zero spatial autocorrelation). When interpreting local autocorrelation measures, it is important to consider both the effect size estimates and the significance level. Since the significance level is calculated for each spot separately, it is recommended to adjust for multiple testing (e.g., Benjamini and Hochberg [benjamini_hochberg1995fdr]). Overall, local Moran’s statistics reveal locations that have similar values to their neighbours. However, as can be seen in Figure 3A-C regions with the highest and most significant local Moran’s value lie in the part of the tissue where expression of gene Nrgn is very low amongst neighbours. Notably, the local Moran’s measure is both dependent on the value of the log-transformed counts and the similarity among neighbours (c.f., Supplementary Figure S2A-B). For the analysis of categorical data, there are metrics such as the join count statistics [daleSpatialAnalysisGuide2014, pp. 141ff.]. In essence, the join count statistic calculates the frequency of categories among neighbours and compares this value with a theoretical distribution or permutations of the labels to get a significance score. This can be used to investigate the interactions between cell types in a lattice since the distance can be specified when constructing the neighbourhood graph.
Challenges and limitations
The choice of the weight matrix poses a challenge in the analysis of lattice data, as it might influence downstream analyses and conclusions. A common choice is to give a non-zero weight only to direct neighbours of a spatial unit (i.e., contiguity-based). Neighbours based on graphs or on distance [pebesmaSpatialDataScience2023, pp. 191 ff.] are also used; the open question is which is best suited for spatial omics data. If one uses contiguity-based neighbours in imaging-based spatial transcriptomics, results ultimately depend heavily on the accuracy of the cell segmentation, which is known to be challenging [stringerCellposeGeneralistAlgorithm2021, greenwaldWholecellSegmentationTissue2022, petukhovCellSegmentationImagingbased2022, architSegmentAnythingMicroscopy2023, wangCellSegmentationImage2019]. If neighbours beyond the adjacent ones are used, the scale (e.g., at which cells interact still) needs to be specified. Figure 3C-D show the difference between the local Moran’s calculation when based on contiguity-based neighbours in D, the 10 nearest neighbours in E or neighbours within a 1000 pixel distance () in F. While the overall differences are small, some cells show different local Moran’s values. Smoothing of the local Moran’s values occurs when more neighbours are considered (Supplementary Figure S2C). Supplementary Figure S2A and B further show that for some cells, no contiguous neighbours were found, which results in a zero estimate of the local Moran’s . Since cells do not function in isolation but form complex anatomical structures including extracellular components, an open question is whether analyses improve when weight matrix construction takes anatomical structures or regions into account. Overall, it remains to be investigated how much the construction of the weight matrix influences downstream analyses in spatial omics data.
Short guide to spatial analysis of biological data
The large variety of options for analysis in spatial statistics can be challenging to navigate. A few fundamental steps should help selecting appropriate exploratory analysis. First, analysts should consider their technology and what options they have (i.e., imaging-based or HTS-based approaches). As seen in Figure 1, the data representation leads to either point pattern or lattice data analysis. If both streams are possible, the decision depends on the research question. In point pattern analysis, the process leading to the locations of individual points is the primary interest. It allows us to make statements such as whether a point process exhibits clustering or if two cell types tend to interact more than randomly expected. In lattice data analysis, however, we are not directly interested in the arrangement of points, but rather locations are important to define the spatial relationship between locations (i.e., via the weight matrix). We study the interaction of features, such as gene expression, given their spatial location. For example, if ligand-receptor dynamics are of interest, the weight matrix can be specified in a way that captures spatial associations of gene expression between direct neighbours or at a specified scale.
Second, the observation scale has to be chosen. Sometimes, this is a technological scale such as multiple FOVs in imaging. Figure 4A shows a single FOV with islets from the human pancreas acquired using IMC [damondMapHumanType2019]. An analysis considering the entire FOV will tell us about the processes that gave rise to the spatial distribution of the endocrine pancreas (islets). As we see in Figure 4B, a Ripley’s -function under a homogeneity assumption indicates that in this observation window, endocrine cells are clustered within pancreatic islets. However, there might also be a relevant biological scale where we subset our observation window to the islets themselves (Figure 4D). In this case, Ripley’s -function indicates a homogeneous distribution of the islets cells within the observation window (Figure 4E). Overall, the correct scale needs to be chosen in accordance with the research question, since the interpretation depends on it.
In point pattern analysis, the analyst needs to decide whether (and how) to control for inhomogeneous intensities, leading back to potential confounding of intensity and interaction [baddeleySpatialPointPatterns2015, pp. 151 ff.]. Analysts can look at smoothed density estimates to inspect potential inhomogeneity prior to analysing the pattern more formally. Furthermore, it depends again on the research question and biological domain knowledge whether the assumption of homogeneity is appropriate. While tissues such as the adipose tissue form regular structures that could be assumed to be homogeneous, other tissues such as stratified epithelial structures appear to be more inhomogeneous. As shown in Figure 4, the homogeneous and inhomogeneous variants of the spatial summaries can lead to different interpretations.
A fundamental choice in lattice data analysis is the construction of the weight matrix. As seen in Figure 3A-B, there are differences in the interpretation of results between different weight matrix choices. The appropriate weight matrix design depends on the research question. For instance, if the feature of interest is a ligand that requires direct cell contact, using contiguous neighbours is a reasonable choice. If the feature of interest is a molecule that diffuses to nearby cells (e.g., paracrine signalling), a nearest neighbour or distance-based approach should be considered for constructing the neighbourhood matrix.
Overall, we advise analysts to consider both scale and homogeneity carefully, keeping the biological context in mind and to document the decisions made and their rationale.
Discussion
In this work, we presented multiple options for the exploratory data analysis of spatial omics data. We introduced the foundations of both point pattern and lattice data analysis and showed the usability of both frameworks in accompanying vignettes.
Fundamentally, point pattern and lattice data analysis are two distinct paths to analysing spatial omics data. Analysts should be aware of the assumptions and interpretations that can be drawn from each approach. In lattice data analysis, we regard the spatial coordinates as fixed and focus on the dependence between neighbours, i.e., weight matrices that reflect these neighbour interactions. In contrast, point pattern analysis concentrates on the underlying mechanisms that generate stochastic spatial patterns (i.e., clustered, regular or spaced arrangements). Therefore, the choice of the analysis framework depends not only on the data modalities, but also on the research question [anselin2024introduction, cressieStatisticsSpatialData1993, baddeleySpatialPointPatterns2015, pebesmaSpatialDataScience2023, daleSpatialAnalysisGuide2014]. Meanwhile, there are dualities between the data modalities from spatially-resolved measurements (c.f., Figure 1). Using cell segmentation, imaging data can be approximated as a points via their centroids. Point locations are stochastic and can therefore be regarded as a realisation of a point process and can be studied using point pattern analysis. On the other hand, we can also view segmentations as an irregular lattice (and aggregate expression per cell). Similarly, HTS-based data with subcellular resolution (e.g., Visium HD [oliveiraCharacterizationImmuneCell2024]) can likewise be interpreted as a regular lattice of (sub)-cellular locations or irregular lattice from segmentations [polanskiBin2cellReconstructsCells2024] or the reconstructed cells can then be approximated by their centroids as points and further analysed using point pattern analysis.
We have focused here on spatial statistics and highlighted two main streams of analysis. Another popular way to represent spatial arrangements is spatial graphs [aliGraphCompassSpatialMetrics2024, windhagerEndendWorkflowMultiplexed2023]. The methods described in the section on lattice data are closely connected to spatial graphs, especially in the construction of a weight matrix. Point pattern analysis concepts are also related to spatial graph via edge rules. For example, Ripley’s can be interpreted as a graph problem with a distance threshold edge rule, as explained in [daleSpatialAnalysisGuide2014, p. 136].
One common challenge in all spatial analyses is the question of the correct scale. In Figure 4, we showed that the scale influences the interpretation of the spatial analysis: what is found to be spatially heterogeneous at a one scale can be locally homogeneous at another [turnerEffectsChangingSpatial1989, daleSpatialAnalysisGuide2014]. Thus, the scale should be defined in accordance with the scientific question.
[…]
There are various other limitations when studying spatially-resolved omics data. On the technical side, imperfect sections can lead to artifacts or non-overlapping FOVs, which can introduce missing data. Furthermore, cells are never perfectly arranged within a histological section, thus cells overlapping in the z-axis are to be expected. In terms of analysis, two-dimensional views neglect processes that occur in three dimensions. For example, in lattice data we might neglect an important contiguous neighbour in an adjacent slice. The same applies to point pattern analysis where we consider only processes in two dimensions even though the underlying biological process likely happens in three dimensions. Both lattice data and point pattern analysis extend conceptually to 3D, however with increased computational complexity. Moreover, in point pattern analysis, we often implicitly assume that patterns are invariant to rotations (see box Assumptions). This assumption is not anatomically accurate for organs and tissues with layered structures, such as the brain. There are other limitations arising from the concept of spatial metrics. Lattice data analysis methods were originally designed for rather low dimensional datasets, while spatial omics often deals with thousands of features (e.g., full transcriptome). Similarly, point pattern methods usually only allow for pairwise comparison of features, which can lead to a high number of cross comparisons e.g., when there are a dozen cell types. Moreover, with newer technologies, the size of the biological sample and resolution of measurements are constantly expanding, which can lead to computational performance issues (e.g., to store cell segmentations in memory to construct weight matrices). Since point pattern based methods only store coordinates, they usually scale better in terms of memory requirements.
In our accompanying vignette, we discuss concepts and assumptions in more detail using biological examples with R and Python. For point pattern analysis, we mainly use R due to the lack of comprehensive Python libraries available for point pattern analysis. Since the R package Voyager is an extensive framework and resource for lattice data, we build on it for the lattice data analysis part of our work. In addition, we provide vignettes for point pattern analysis to complement existing discussions [zormpasMappingTranscriptomeRealizing2023, mosesVoyagerExploratorySinglecell2023].
Outlook
The spatial metrics as described in point pattern and lattice data analysis offer interesting ways to quantify spatial patterns in molecular biology. There are still open research questions including the correct application (e.g., scale and homogeneity) of methods to a given biological context, the high dimensional nature of spatial omics data and the effect of preprocessing steps that come upstream of the application of spatial statistics. Furthermore, comparisons of spatial metrics across multiple samples and conditions and parametric spatial models offer interesting options for future work.
Code and data availability
Code and instructions to retrieve data to reproduce figures and results can be found under https://github.com/robinsonlabuzh/pasta-manuscript, code licensed under the GPLv3 terms. The vignette and source code can be accessed via https://robinsonlabuzh.github.io/pasta licensed under CC-BY-4.0 terms for the text and GPLv3 license for any code.
Acknowledgments
We thank all members of the Robinsonlab for constructive feedback on the manuscript and vignette. In particular, we thank Peiying Cai, Alice Driessen, Reto Gerber, Pierre-Luc Germain, Maruša Kodermann, Vladyslav Korobeynyk, Siyuan Luo, Giulia Moro, Emanuel Sonder, Jiayi Wang and David Wissel for their careful reading and input to the vignette. The Python examples in the vignette were developed as part of the SpaceHack 2024 event. We thank the SpaceHack organisers and participants.
Author contributions
ME: Conceptualisation, Methodology, Software (Vignette), Analysis, Visualisation, Writing - Original Draft; SG: Conceptualisation, Methodology, Software (Vignette), Analysis, Visualisation, Writing - Original Draft; HLC: Conceptualisation, Vignette, Writing - Review & Editing; IM: Methodology, Writing - Review & Editing; MK, Software (Vignette), Writing - Review & Editing; RF: Methodology, Writing - Review & Editing; MDR: Conceptualisation, Methodology, Writing - Original Draft, Supervision, Funding acquisition.
Funding
This work was supported by Swiss National Science Foundation (SNSF) project grant 310030_204869 to MDR. MDR acknowledges support from the University Research Priority Program Evolution in Action at the University of Zurich. HLC acknowledges support by SNSF grant number 222136.
Conflict of interest
We declare no conflict of interest.
Supplementary Figures

### v3

@@
-pasta: Pattern Analysis for Spatial Omics Data
+Harnessing the Potential of Spatial Statistics for Spatial Omics Data with pasta
 Abstract
-Spatial omics assays allow for the molecular characterisation of cells in their spatial context.
-Notably, the two main technological streams, imaging-based and high-throughput sequencing-based, can give rise to very different data modalities.
-The characteristics of the two data types are well known in adjacent fields such as spatial statistics as point patterns and lattice data, and there is a wide range of tools available.
-This paper discusses the application of spatial statistics to spatially-resolved omics data and in particular, discusses various advantages, challenges, and nuances.
-This work is accompanied by a vignette, pasta (https://robinsonlabuzh.github.io/pasta), that showcases the usefulness of spatial statistics in biology using several R and Python packages.
+Spatial omics allow for the molecular characterization of cells in their spatial context.
+Notably, the two main technological streams, imaging-based and high-throughput sequencing-based, give rise to very different data modalities.
+The characteristics of the two data types are well known in spatial statistics as point patterns and lattice data.
+In this perspective, we show the versatility of spatial statistics to quantify biological phenomena from local gene expression to tissue organization.
+As an example, we describe how to use exploratory metrics to address scientific questions in breast cancer, including cellular co-localization and gene co-expression analysis.
+We discuss technical concepts like window sampling, homogeneity and weight matrix construction and show their importance.
+We also provide pasta (https://robinsonlabuzh.github.io/pasta), an extensive analysis vignette for spatial statistics both using R and Python packages with further biology-driven applications.
 Introduction
-Molecular profiling of cells in organs or tissues can be accomplished either in bulk or at the single-cell level.
-However, tissue must be dissociated, which may select against certain cell types and results in the loss of the spatial organization of the cells.
-This limitation is overcome with spatial profiling techniques, which can range from spatial proteomics (e.g., IMC, MIBI-TOF, 4i, CODEX, mIF [giesenHighlyMultiplexedImaging2014, kerenMIBITOFMultiplexedImaging2019, gutMultiplexedProteinMaps2018, blackCODEXMultiplexedTissue2021, tsujikawaQuantitativeMultiplexImmunohistochemistry2017]), to spatial transcriptomics assays based on fluorescence in situ hybridisation (e.g., FISH, MERFISH, CosMx, Xenium [langer-saferImmunologicalMethodMapping1982, chenSpatiallyResolvedHighly2015, janesickHighResolutionMapping2023]) and based on sequencing (e.g., 10x Visium and Visium HD, Slide-seq V1&V2 [stahlVisualizationAnalysisGene2016, oliveiraCharacterizationImmuneCell2024, rodriquesSlideseqScalableTechnology2019, stickelsHighlySensitiveSpatial2021]) to even spatial epigenomics [luSpatiallyResolvedEpigenomic2022].
-Moreover, the combination of individual methods enabled spatial multiomic research [vandereykenMethodsApplicationsSinglecell2023].
+Molecular profiling at spatial resolution
+Molecular profiling of cells in bulk or at the single-cell level can be accomplished by dissociating organs or tissues.
+Dissociation may select against certain cell types [denisenkoSystematicAssessmentTissue2020, dingSystematicComparisonSinglecell2020] and results in the loss of the spatial organization.
+This limitation is attenuated with spatial profiling techniques, which preserve the cellular spatial organization and range from spatial proteomics [giesenHighlyMultiplexedImaging2014, kerenMIBITOFMultiplexedImaging2019, gutMultiplexedProteinMaps2018, blackCODEXMultiplexedTissue2021, tsujikawaQuantitativeMultiplexImmunohistochemistry2017], to spatial transcriptomics based on fluorescent in situ hybridisation [langer-saferImmunologicalMethodMapping1982, chenSpatiallyResolvedHighly2015, janesickHighResolutionMapping2023] or on sequencing [stahlVisualizationAnalysisGene2016, oliveiraCharacterizationImmuneCell2024, rodriquesSlideseqScalableTechnology2019, stickelsHighlySensitiveSpatial2021, chenSpatiotemporalTranscriptomicAtlas2022], to spatial epigenomics [luSpatiallyResolvedEpigenomic2022] and spatial metabolomics [angelMatrixAssistedLaserDesorption2013, zhangMassSpectrometryImaging2023], or multi-omics [vandereykenMethodsApplicationsSinglecell2023].
 The technologies and their application in biological research have been the topic of various detailed reviews [aspSpatiallyResolvedTranscriptomes2020, moffittEmergingLandscapeSpatial2022, raoExploringTissueArchitecture2021, mosesMuseumSpatialTranscriptomics2022, pallaSpatialComponentsMolecular2022, bressanDawnSpatialOmics2023].
-In this work, we focus on the concepts and application of exploratory statistical approaches for spatial data with a focus on downstream steps; our overview also centres around key differences between data modalities that result from current spatial profiling technologies.
-In this regard, we complement the ongoing discussion [zormpasMappingTranscriptomeRealizing2023, mosesVoyagerExploratorySinglecell2023] about the challenges in applying spatial statistics to spatial (transcript)omics data.
-Most spatial omics assays can be classified as either high-throughput sequencing (HTS) based or imaging-based.
-In HTS-based approaches, positional information is recorded using a predetermined array of spots or beads.
+Spatial data modalities have been present in other fields for decades.
+They originated in mining and forestry and have expanded to other fields such as geography or ecology, and have been approached by spatial statistics methods [zuurAnalysingEcologicalData2007, daleSpatialAnalysisGuide2014, baddeleySpatialPointPatterns2015, pebesmaSpatialDataScience2023].
+In this work, we focus on the concepts and application of exploratory statistical approaches for molecular spatial data; our overview also centres on key differences between data modalities that result from current spatial profiling technologies and the different types of questions that can be answered with spatial statistics.
+In this regard, we complement the ongoing discussion on the challenges in applying spatial statistics to spatial (transcript)omics data [zormpasMappingTranscriptomeRealizing2023, mosesVoyagerExploratorySinglecell2023, gonzalez2025analyzing].
+Data modality is relevant to spatial data analysis
+Most spatial omics assays can be classified as either high-throughput sequencing (HTS)-based or imaging-based.
+In HTS-based approaches, positional information is recorded by barcoding a predetermined array of spots or beads.
 Imaging-based approaches, however, either target the molecules of interest with fluorescent probes, ablate regions stained with a cocktail of antibodies via metal tag readouts, or amplify and sequence target sequences in situ.
 Several technologies are emerging, but the main trade-offs stem from the resolution, number of features, and sensitivity of the readout [raoExploringTissueArchitecture2021, parkSpatialOmicsTechnologies2022].
-For example, HTS-based approaches capture the entire transcriptome (i.e., untargeted), but come at a resolution determined by the spot size.
-On the other hand, imaging-based approaches typically have lower depth of information (i.e., targeted) but show a higher resolution in comparison to HTS-based approaches.
-Note also that the landscape of spatially-resolved molecular technologies is rapidly changing, mostly in terms of higher resolution for HTS-based methods and comprehensive marker panels for imaging-based approaches (Figure 1, Technology) [moffittEmergingLandscapeSpatial2022, raoExploringTissueArchitecture2021, rademacherComparisonSpatialTranscriptomics2024].
-In terms of data analysis, the technological streams are quite distinct.
-HTS-based approaches collect data along regularly-spaced spots or beads resulting in a so-called regular lattice.
-In contrast, imaging-based approaches measure features at exact locations that can be assumed to have originated from a stochastic process known as a point process [raoExploringTissueArchitecture2021].
-Thus, we will distinguish between lattice-based and point pattern-based spatial omic data.
-While the focus is often on the technology (c.f., imaging-based versus HTS-based), we argue that the distinction should lie at the data representation level [raoExploringTissueArchitecture2021].
-This is important, since there are technologies that can be represented as both point patterns and lattice data, depending on the resolution and data processing.
-For example, imaging-based technologies that detect transcripts at subcellular resolution result in data where transcript locations can be represented as a point pattern, while for the same dataset, segmented cells can be represented either as an irregular lattice, or as a point pattern via their cell centroids.
-In general, if events of a point pattern are aggregated in specified regions (e.g., transcripts per cell from segmentation), we end up with lattice data (Figure 1, Data modality) [daleSpatialAnalysisGuide2014, pebesmaSpatialDataScience2023].
-Since we can represent some spatial omic datasets as either lattice or point pattern, it is important to further understand the assumptions and analysis strategies of such data modalities [cressieStatisticsSpatialData1993].
-On the one hand, spatial point pattern analysis assumes that the point locations were generated by a stochastic process, a so-called event-based view of the data; the goal is to study the properties of this process [baddeleySpatialPointPatterns2015].
-On the other hand, one could assume that the locations are fixed and known at the time of sampling and study the associated features at each location via an observation-based view of the data, while recognising that the observations in the lattice are not independent due to the spatial structure [anselinWhatSpecialSpatial1989].
-Spatial data modalities have been present in other fields for decades, such as geography, and in particular, the field of spatial statistics offers a large analysis toolbox.
-For example, spatial omics data collected across a lattice can be analysed by exploratory tools, such as Local Moran’s and the bivariate Lee’s statistic.
-Imaging data generated by a stochastic process can be input for methods of point pattern analysis, including summaries such as Ripley’s or empty-space functions [zuurAnalysingEcologicalData2007, daleSpatialAnalysisGuide2014, baddeleySpatialPointPatterns2015, pebesmaSpatialDataScience2023] (Figure 1, Analysis).
-There are already several tools for omics analysis that employ spatial statistics approaches.
-Amongst others, Voyager is a framework and a collection of use cases for lattice data, written in both R and Python [mosesVoyagerExploratorySinglecell2023]; it provides data structures and functionalities to compute spatial statistics in molecular biology.
-Voyager comes with comprehensive vignettes for several spatial profiling technologies.
-The methodological focus of their vignettes is on lattice data analysis and the spdep geospatial package.
-Since Voyager offers efficient implementations for lattice data analysis in R, we build on their framework in our vignettes for the lattice data component.
-Other methods written in R include: SPIAT, which contains various tools for spatial analysis, such as: neighbourhood analysis, local metrics and heterogeneity scores etc.
-[fengSpatialAnalysisSPIAT2023]; scFeatures, a toolbox comprising lattice data analysis functionality such as Moran’s and point pattern metrics as the -function [caoScFeaturesMultiviewRepresentations2022]; spicyR, which uses a compressed version of the global -function for cross-sample comparison [caneteSpicyRSpatialAnalysis2022]; lisaClust, a spatial domain detection method via LISA (local indicators of spatial association) curves [patrickSpatialAnalysisHighly2023]; mxfda, that uses point pattern summary functions for survival analysis [wrobelMxfdaComprehensiveToolkit2024]; spatialDM uses global and local bivariate Moran’s to score colocalisation of ligand-receptor pairs [liSpatialDMRapidIdentification2023]; and MERINGUE, which uses nearest-neighbour autocorrelation for spatially variable gene selection [millerCharacterizingSpatialGene2021].
-Methods in Python include squidpy, a package with various spatial statistics tools for both point pattern and lattice data [pallaSquidpyScalableFramework2022] and PySAL a general Python library for spatial analysis [reyPySALPythonLibrary2010].
-For a detailed view on general methods development in the field of spatial transcriptomics, we refer readers to the online Supplementary table of the museum of spatial transcriptomics paper [mosesMuseumSpatialTranscriptomics2022].
-Here, we will explore the application of spatial statistics to spatially-resolved omics data guided by the two main streams of data, point pattern-based and lattice-based methods.
-We will compare the two streams and show their strengths and limitations.
-Finally, we give an outlook on the challenges and research gaps in the field of spatial data analysis for omics data.
-Furthermore, we provide a vignette (https://robinsonlabuzh.github.io/pasta) that showcases analysis of data from multiple technologies, where concepts and assumptions are discussed in detail, with inline R and Python code.
-Overall, our resource Pattern Analysis for SpaTial omics datA (pasta) will highlight the usefulness and transferability of existing exploratory spatial statistics approaches in the context of spatial tissue profiling.
-Results
-Point Pattern Analysis
-Definitions
-Most imaging-based technologies give a high-resolution readout of either subcellular compartments (e.g., laser ablation pixels in IMC [giesenHighlyMultiplexedImaging2014]) or even individual transcripts (e.g., FISH [feminoVisualizationSingleRNA1998]).
-In molecule-level technologies, the expression of features (transcript quantifications, ion counts, fluorescent intensities) is recorded where they occur [aspSpatiallyResolvedTranscriptomes2020].
-Often, we are interested in the distribution and composition of cell types.
-In order to annotate cells, we require (good) cell segmentation boundaries (an active area of research not discussed here; see [wangCellSegmentationImage2019, stringerCellposeGeneralistAlgorithm2021, greenwaldWholecellSegmentationTissue2022, petukhovCellSegmentationImagingbased2022, architSegmentAnythingMicroscopy2023]), allowing transcripts to be allocated to a given cell [fuBIDCellBiologicallyinformedSelfsupervised2024], and then cell type inference.
-Since the locations of the annotated cells are generated by a stochastic biological process, we can approximate e.g., the cell centroids as points and analyse their location as point patterns.
-In the analysis of these point patterns, the goal is to make inferences on the point process that generated the data, not on the patterns themselves [baddeleySpatialPointPatterns2015, p.
+For example, HTS-based approaches aim to profile the entire transcriptome (i.e., untargeted), but come at a resolution determined by the spot size.
+On the other hand, imaging-based approaches typically describe a lower number of features (i.e., targeted) but show a higher resolution.
+Note also that the landscape of spatially-resolved molecular technologies is rapidly changing, mostly in the direction of higher resolution for HTS-based methods and higher-plex marker panels for imaging-based approaches (Figure 1A; Supplementary Table S1) [moffittEmergingLandscapeSpatial2022, raoExploringTissueArchitecture2021, rademacherComparisonSpatialTranscriptomics2024].
+In terms of data analysis, HTS- and imaging-based outputs are quite distinct.
+HTS-based approaches collect data along regularly spaced spots or beads resulting in a so-called regular lattice.
+In contrast, imaging-based approaches measure features at exact locations that can be assumed to have originated from a random (stochastic)process known as a point process [raoExploringTissueArchitecture2021], leading to so-called point patterns.
+Although a random process might not be the correct assumption in biological development, the representation of point patterns can still be useful when correctly interpreted as we will show later.
+Thus, we will distinguish between lattice-based and point pattern-based spatial omics data.
+While the classification is often on the technology (c.f., imaging-based versus HTS-based), we argue that the distinction should lie at the data representation level [raoExploringTissueArchitecture2021].
+This is important, since there are technologies that can be represented as both point patterns and lattice data depending on the resolution and data processing.
+For example, imaging-based data can be processed to describe the location of single transcripts at subcellular resolution (point pattern) while also identifying individual cells as either an irregular lattice when using the cell segmentation as the lattice structure or as a point pattern when considering cell centroids.
+In general, if events of a point pattern are aggregated into specified regions (e.g., transcripts per cell from segmentation), we end up with lattice data (Figure 1B) [daleSpatialAnalysisGuide2014, pebesmaSpatialDataScience2023, cressieStatisticsSpatialData1993].
+Lattice and point pattern analysis represent two overlapping streams of spatial statistics
+Since we can represent some spatial omics datasets as either lattice or point pattern, it is important to further understand the assumptions underlying common data analysis strategies.
+On the one hand, point pattern analysis assumes that the point locations were generated by a random process, where each point is an event itself, a so-called event-based view of the data; its goal is to study the data generating process [baddeleySpatialPointPatterns2015].For example, if we are interested in the distribution of one cell type in an image, we would use point pattern methods to quantify this arrangement.
+On the other hand, one could assume that the locations are fixed and known at the time of sampling and study the associated features at each location via an observation-based view of the data, while recognizing that the observations in the lattice are not independent measurements due to their spatial dependence [anselinWhatSpecialSpatial1989].
+For example, in spot-based technologies, locations are predefined.
+If we want to study a gene expression pattern in such an assay, we should account for the spatial dependence between spots (Figure 1C).
+In addition, characterizing the marks (variable) of interest is also important.
+We consider two broad types of marks associated with spatial locations: categorical (qualitative, e.g.
+cell types or spatial domains) and numerical (quantitative, e.g.
+gene expression).
+In terms of dimensionality, marks can be analyzed one- (univariate), two- (bivariate) or three or more at a time (multivariate analysis).
+Figure 1C summarizes options to analyze each combination of the type of mark and its dimensionality.
+Defining characteristics of marks is relevant to choose appropriate metrics for the biological question in mind; e.g., infiltration of cell type A into a region dominated by cell type B is a question involving a categorical bivariate measure (two cell types involved) and can be analyzed as a type of co-localization analysis between type A and B cells.
+Notably, there are more methods for the analysis of categorical marks with point patterns.
+For numerical marks, more options exist in lattice data analysis (Figure 1).
+This should serve as an indication of which stream can be useful, if the data modality from the experimental assay allows for both point pattern and lattice data analysis.
+![[Uncaptioned image]](x1.png)
+(Figure legend on next page.)
+Here, we will explore the application of spatial statistics to spatially-resolved omics data guided by concrete biological questions (Figure 1).
+We will further focus on the two main streams of data analysis, point pattern-based and lattice-based methods, describing their strengths and limitations.
+Finally, we will give an outlook on the challenges and research gaps in the field of spatial data analysis for omics data.
+Biological applications of spatial statistics for spatial omics
+Spatial autocorrelation reveals triple positive regions in breast tumours
+We re-analyzed an imaging-based human breast cancer dataset [janesickHighResolutionMapping2023], where the cell-level data can be considered as both a point pattern or lattice data (Figure 1).
+Figure 2A shows the H&E image of the breast cancer section.
+As the first feature of interest, we consider the measured gene expression, which is a numerical, continuous (cell-level) “mark”.
+As we are interested in relationships within the local neighbourhood of the cells rather than effects at specific locations or distances, lattice data is the appropriate analysis method (Figure 1).
+In their publication, Janesick et al.
+[janesickHighResolutionMapping2023] investigate the expression of three clinically relevant hormone receptor genes: Estrogen receptor (ESR1), progesterone receptor (PGR) and human epidermal growth factor receptor 2 (ERBB2).
+The presence or absence of these receptors (indicated with +/-) can inform treatment strategies [janesickHighResolutionMapping2023].
+The expression pattern of each receptor is quite distinct, with ERBB2 being more highly expressed compared to both ESR1 and PGR (Figure S1A-C).
+Lattice data analysis enables us to use a locally-scaled measure of expression to distinguish high and low expression regions.
+Since we are interested in localized hormone receptor-positive regions, we define the neighbourhood as the six nearest neighbours around each cell.
+Univariate spatial autocorrelation allows us to identify regions where the expression of the hormone receptor genes is correlated among cells in their neighbourhood.
+Since autocorrelation measures such as local Moran’s [moranNotesContinuousStochastic1950, anselinLocalIndicatorsSpatial1995] are non-directional (e.g., local autocorrelation can be both high in regions where expression among neighbours is high or low as the focus in on similarity; for more details, see section “Spatial autocorrelation is a measure of similarity”), we rely on Moran’s scatter plot to identify the type of interaction [anselin2019moran].
+In regions with “high-high” interaction, the expression among neighbours is both similar and high, which can be interpreted as a region positive for one of the receptors (Supplementary Figure S1D-F).
+Figure 2B shows the union of the individual positive univariate classifications of the three receptor genes and indicates single, double and triple positive receptor regions.
+One DCIS region (indicated by arrow) is lined with predominantly triple positive cells, but some local heterogeneity with double and single positive cells is apparent.
+We also applied bi- or multivariate spatial correlation measures to classify the different regions.
+Similar to spatial autocorrelation, bivariate autocorrelation measures the similarity of expression of two genes in a local neighbourhood.
+Supplementary Figure S2A-C show bivariate local Lee’s [leeDevelopingBivariateSpatial2001] correlation measures for the pairs of receptors ERBB2-ESR1, ERBB2-PGR and ESR1-PGR, respectively; regions with high values indicate regions where expression of both receptor genes is similar (“high-high” or “low-low”).
+The different scales of the local Lee’s L values are the result of the different expression scales (c.f., Supplementary Figure S1A-C).
+A comparison of gene expression values and the local Lee’s value allows us to identify double positive regions.
+Multivariate spatial correlation metrics extend the concept of spatial correlation to more than two genes.
+The main difficulty is finding an appropriate metric to measure similarity of more than two marks.
+In multivariate Geary’s , the respective metric is the weighted average of the squared differences of the expression vectors of the three genes.
+Supplementary Figure S2D-F shows an analysis of multivariate local Geary’s [anselinLocalIndicatorMultivariate2019] for the three hormone receptors.
+Cells with a high local Geary’s value exhibit a gene expression profile that is highly dissimilar (squared difference is high) to their local neighbours and indicates local heterogeneity.
+Interestingly, the triple positive region identified using Moran’s scatter plot shows high multivariate Geary’s values.
+At first, this seems counterfactual to what we have identified using the Moran’s scatter plot.
+A closer look reveals that the region is very heterogeneous with a lot of single and double-positive regions with varying scales of expression of the individual genes.
+Generally, there is a difficulty in interpreting multivariate autocorrelation measures as they are not simply the sum of the respective univariate counterparts and should therefore be mainly used to identify interesting regions (c.f., Anselin et al.
+[anselinLocalIndicatorMultivariate2019]).
+In summary, we have investigated the expression patterns of the three clinically-relevant receptor genes, quantifying single-, double- and triple-positive receptor regions in the tumour using lattice data analysis.
+Spatial statistics was able to recapitulate the original findings in Janesick et al.
+[janesickHighResolutionMapping2023] (Figure 5), and in addition, spatial exploratory statistics provide a quantification of the strength of the various spatial correlations as well as an indication of the significance of the effects.
+Point pattern analysis quantifies different degrees of tumour invasion into ductal carcinomas
+Another analysis to consider is the exploration of the spatial distribution of cell types in the Janesick et al.
+[janesickHighResolutionMapping2023] dataset.
+Since cell types are categorical features, point pattern analysis is a powerful choice [baddeleySpatialPointPatterns2015] (Figure 1).
+Here we consider the DCIS 1, DCIS 2 and invasive tumour cell types and aggregate all other cell types into the “Other” category (Figure 3A).
+In this example, we calculate Besag’s cross function to quantify the co-localization of the three cell types in space [besag1977contribution, ripleySecondOrderAnalysisStationary1976].
+Note that the observation window was restricted to the region where most of the cells were found to account for the underlying inhomogeneous distribution of cells (c.f., Figure 3).
+There are other options to account for this inhomogeneity, which are compared in Figure S3.
+In general, Besag’s calculates the average number of neighbours of a certain cell type as a function of the radius.
+This can used to estimate the magnitude of co-occurrence relative to a scenario where all cells are randomly distributed in the observation window.
+In Besag’s , this corresponds to a value of 0 across all radii.
+Positive deviation indicates more cells are observed than expected by chance, implying clustering; negative deviation indicates less cells observed than expected by chance, implying spacing of the cells.
+First, we consider univariate co-localization, i.e., whether cells are found more often than expected in the vicinity of other cells from the same type.
+Figure 3B shows Besag’s curve plots for each pair of cell types.
+We note that all cell types exhibit curves above the (dashed black) reference line, which means that we observe, on average, more cells of the same type within a certain radius than expected by chance.
+This represents a higher degree of clustering than a fully random, or completely spatial random (CSR), point pattern (see Technical description below; c.f., box “Importance of scale and uniformity in point pattern analysis”).
+Thus, all three cell types have a propensity higher than expected by chance to occur in the neighbourhood of their own type (Figure 3A) and this clustering is strong across all scales (or radii, ).
+Besag’s can quantify bivariate interactions, i.e., the co-localization of a pair of cell types, as shown in Figure 3B.
+We note that the combination of DCIS 1 and DCIS 2 is below the dashed CSR line.
+This means that we observe less cells of the other type within a specific radius than expected by chance.
+This can be interpreted such that there is, on average, spacing between the respective cell types.
+Of interest is the co-localization of the two DCIS subtypes and invasive tumour cells: DCIS 1 cells and invasive tumour cells are strongly spaced, thus clearly co-occur less frequently than expected at random.
+On the other hand, DCIS 2 cells and invasive tumour cells are close to the CRS line which means they occur as often as expected by chance, an indication of mixing of the two cells types in space.
+In summary, we have recapitulated the main findings on DCIS cells presented in Janesick et al., Figure 4, using point pattern analysis [janesickHighResolutionMapping2023].
+DCIS subtypes show a different degree of co-localization with invasive tumour cells, with DCIS 1 regions showing no co-localization with invasive tumour while DCIS 2 exhibits a degree of co-localization with invasive tumour cells that can be expected by chance.
+Using spatial point pattern analysis, we have systematically quantified the co-localization of three cell types in breast cancer.
+Further analyses using spatial statistics
+In addition to the analyses described above, spatial statistics could be used for various other analyses.
+Univariate global spatial autocorrelation methods such as Moran’s can be used to identify the strength of spatial correlation across all genes in the panel.
+This information can be useful to compare spatial variability of different genes within one sample and the specified neighbourhood.
+In addition, by constructing a spatial weight matrix based on contiguous neighbours (i.e., cells in physical contact with each other), we can identify pairs of ligand-receptor genes via quantification of their bivariate spatial correlation [liSpatialDMRapidIdentification2023].
+Alternatively, adjusting the weight matrix to neighbours based on a certain distance, we can investigate paracrine signalling up to the distance specified.
+Investigating the spatial arrangement of cells with point pattern analysis, we could further use mark correlation functions to quantify the scale of spatial correlation of the expression of a gene of interest.
+In addition, we could profile the co-localization of cell types using nearest-neighbour approaches such as the nearest-neighbour distance and orientation.
+These metrics would provide us with a more quantitative readout of the spatial organization than we can identify by eye.
+“Dot” (or “-to-any”) functions would allow us to make statements about how one cell type is localized with respect to a set of other cell types.
+Moreover, we could also focus our analyses to regions of interest such as the single-, double- and triple-positive tumour regions to make statements about localized gene expression and the arrangement of cell types within these regions.
+The observation scale is very important in spatial analysis as it will impact the interpretation of spatial relationship, as we will further explain below.
+![[Uncaptioned image]](x3.png)
+(Figure legend on next page.)
+Technical considerations
+Inference about the point process
+In the analysis of point patterns, the goal is to make inferences on the point process that generated the data, not on the patterns themselves [baddeleySpatialPointPatterns2015, p.
 127].
-Points in a point pattern can also carry a mark, which is a (univariate or multivariate) variable associated with a point [baddeleySpatialPointPatterns2015, pp.
-147, 563].
-In cell biology, we could imagine both categorical (e.g., cell type or spatial domain labels) or continuous marks (e.g., gene expression).
-When a point pattern with many cell types is considered, there are two approaches to formulate the setup.
-In the first variant, we consider cells in a tissue to depend on each other so their distribution is due to one overarching biological process; this is referred to as a multitype view.
-In contrast, when we consider the patterns to be created by point processes, we assume that these processes can be individually analysed; this represents the multivariable viewpoint (change in terminology in comparison to [baddeleySpatialPointPatterns2015]).
-The processes do not need to be independent, but the independence of components (and functions) is often used as a null model for multivariable analysis (see box Assumptions) [baddeleySpatialPointPatterns2015, pp.
-565ff.].
-Assumptions
+When considering a point pattern with many cell types, there are two possible formulations.
+In the first variant, we assume that cells in a section depend on each other so their distribution is due to one overarching biological process; this is referred to as a multitype view.
+In contrast, when we consider the patterns are created by point processes, we assume that these processes can be individually analyzed; this represents the multivariable viewpoint (change in terminology in comparison to [baddeleySpatialPointPatterns2015]).
+Generally, it is not easy to define whether the entire pattern of several cell types should be represented as a single point process (multitype analysis).
+Notably, analyzing cell types individually and considering them as separate point processes might underestimate dependencies among the cells (multivariable analysis) [baddeleySpatialPointPatterns2015, pp.
+565 ff.].
+Point pattern analysis quantifies spatial interactions across scales
+A central idea in point pattern analysis is that of an -neighbourhood.
+The -neighbourhood describes the spatial dependence in a disk with radius .
+In such a neighbourhood, different aspects can be quantified, leading to different dependent variables per function.
+The expected number without any spatial interaction (CSR or Poisson process) can be calculated in the -neighbourhood.
+Positive deviations from the CSR value indicate clustering whereas negative deviations indicate spacing [baddeleySpatialPointPatterns2015, p.
+204].
+Therefore, point pattern metrics are functions that quantify spatial dependence across scales.
+The two most well-known classes in point pattern analysis are correlation and spacing functions.
+The two approaches are highly complementary and there is a duality between them [baddeleySpatialPointPatterns2015, p.
+255].
+Correlation is usually assessed with Ripley’s function or its variance-stabilized version, Besag’s .
+Here, the number of points falling in an -neighbourhood is quantified.
+These raw counts are corrected for varying intensities as well as edge effects (see Box Importance of scale and uniformity in point pattern analysis) [ripleyFindingEdgePoisson1977, besag1977contribution][baddeleySpatialPointPatterns2015, p.
+203 ff.].
+Spacing can be assessed with the nearest-neighbour function .
+Here, the probability of finding the nearest-neighbour of a point in an -neighbourhood is calculated [baddeleySpatialPointPatterns2015, p.
+261 ff.].
+Both spacing and correlation quantify different aspects of a point pattern and a spatial analysis should consider both complementary options [baddeleySpatialPointPatterns2015, p.
+295 ].
+Window sampling and small world impact representative sample collection
 Many imaging platforms measure patterns in a field of view (FOV) and collect several such FOVs per sample.
-In some technologies, there are no gaps between the individual FOVs and they can be assembled into one big image.
-If there are gaps, consecutive FOVs can either be stitched together (which can be computationally complex) or analysed individually [aspSpatiallyResolvedTranscriptomes2020, mohammadiFastRobustFeaturebased2024].
-When analysing individual FOVs, we do not observe the entire pattern but rather a subset of the bigger pattern.
-This is called window sampling because the window is just a sample of the bigger point pattern.
+In some technologies, there are no gaps between the individual FOVs so they can be assembled into one image.
+If there are gaps, consecutive FOVs can either be stitched together (which can be computationally complex) or analyzed individually [aspSpatiallyResolvedTranscriptomes2020, mohammadiFastRobustFeaturebased2024].
+When analyzing individual FOVs, not the entire pattern but rather a subset of the full pattern is observed: this is so-called window sampling because the window is a sample of the full point pattern.
 A related concept is the small world model, which describes that points can only be observed in a finite world and not beyond these boundaries [baddeleySpatialPointPatterns2015, pp.
-143ff.].
-Often, the distinction between the window sampling and small world model concepts is not clear.
+143ff., p.
+560].
+Often, the distinction between the window sampling and small world model concepts is not clear in histological sections.
 An example is the arrangement of the epithelium, which can can be imaged with several FOVs (window sampling) but there are no cells expected outside of the epithelial boundary (small world) [garrido-trigoMacrophageNeutrophilHeterogeneity2023].
-Therefore, in cell biology, we often encounter a mixture of these two concepts.
+Therefore, in spatial omics, we often encounter a mixture of these two scenarios.Obtaining a representative sample of a point pattern is also a challenge.
+In spatial omics, we are often provided with FOVs that an experimentalist has selected from a much larger region or that are the result of the data acquisition (e.g., FOVs in CosMx [heHighplexImagingRNA2022]).
+Such a sample might have different characteristics compared to the point pattern of the larger region, as FOVs are sometimes selected based on morphological properties (e.g., H&E staining).
+In cases where we have a small world scenario, we would want to limit our observation window to the full region where the points can occur (c.f., Supplementary Figure S3).
+Importance of scale and uniformity in point pattern analysis
 Apart from the window of measurement, we need to make assumptions on the statistical properties of the point pattern, including, most importantly, whether points can be considered homogeneous or not.
-Homogeneity (see Box Assumptions) assumes that the number of points in a given region is proportional to its area , for arbitrary ; i.e., homogeneity refers to a uniform intensity of points across the window of measurement.
-If that is not the case, the point process is said to be inhomogeneous [baddeleySpatialPointPatterns2015, p.132-133].
-This difference has important implications for the analysis of a process.
-For example, when adjusting for varying local intensities, the process could be found to be inhomogeneous rather than genuinely clustered.
-This is called the confounding between intensity and interaction (see box Assumptions) [baddeleySpatialPointPatterns2015, pp.151-152].
-Options for analysis
-Point pattern analysis offers a broad spectrum of options.
-Most often, we compare patterns in 2D (e.g., a single histological slice), but the methods are applicable with some adjustments to 3D patterns (e.g., confocal microscopy volumes) [baddeleySpatialPointPatterns2015, p.
-651], with some increased computational cost [baddeleyAnalysisThreeDimensionalPoint1993].
-As mentioned above, points can carry marks; the nature of these marks determines what options are available to the analyst.
-For example, many point pattern methods are designed for categorical data, such as cell types or spatial domain labels [baddeleySpatialPointPatterns2015, p.147] and can be adapted to the analysis of continuous marks [baddeleySpatialPointPatterns2015, pp.644 ff.].
-Model-based approaches for spatial data might also be of interest.
-This is an entire subfield on its own with many options available [baddeleySpatialPointPatterns2015, pp.297 ff.][pebesmaSpatialDataScience2023].
-When comparing categorical marks, the analyst must decide whether to compare a single type (within-type comparison) or between a pair of types (cross-type comparison).
-For example, a within-type comparison could use the pattern of epithelial cells in a microscopy image; the correlation, spacing or distances between epithelial cells could be summarised.
-In contrast, a cross-type analysis could consider the correlation between epithelial cells and B cells [baddeleySpatialPointPatterns2015, pp.
-561ff.].
-Examples
-A well-known point pattern summary is Ripley’s -function, defined as the expected number of points falling in a specified radius around an arbitrary point, averaged over all points in the observation window (e.g., FOV) (potentially adjusted for edge effects; see box Assumptions).
-Ripley’s -function thus quantifies the correlation structure of a point pattern, with within-type and cross-type versions.
-A point pattern where points have no preference for location shows complete spatial randomness (CSR).
-This means that the number of points that fall in an arbitrary region follow a Poisson distribution with a rate parameter that is proportional to the size of the region.
-A clustered process would have values of the -function larger than the Poisson process at that specific scale; values of smaller than the Poisson process indicate a self-inhibiting process (i.e., where points repel each other) [ripleySecondOrderAnalysisStationary1976, ripleyModellingSpatialPatterns1977][baddeleySpatialPointPatterns2015, pp.203 ff.].
-Ripley’s -function can be variance stabilised by taking the square root, known as Besag’s -function [besag1977contribution].
-It is a global estimate that measures correlation across an entire observation window and can be used instead of the standard -function [baddeleySpatialPointPatterns2015, p.402], [baddeleySpatstatPackageAnalyzing2005].
-Figure 2A-C show point patterns of mature oligodendrocytes over three different serial sections of the same mouse hypothalamic preoptic region (the numbers , , indicate the -axis position of the slices: one slice every 50 m along the anterior-posterior axis [moffittMolecularSpatialFunctional2018]).
-Figure 2D shows the homogeneous global -function, which quantifies the correlation within each slice and the respective -function for a Poisson process (dashed line).
-The homogeneous -function indicates strong clustering, even for the slice, however, the inhomogeneous -function, which makes a local adjustment for intensity, indicates a level of homogeneity very similar to a (completely spatially random) Poisson process (Supp.
-Figure S1A-C vs.
-D-F).
-This is an example of intensity vs.
-interaction confounding (see box Assumptions).
-If slice should therefore be interpreted as clustered, depends on the assumption of homogeneity for that slice.
-In addition, there are differences between the inhomogeneity corrections in Figure S3; the corrections are discussed in more detail in the vignette.
-Local versions of the various summary functions exist.
-The LISA (local indicators of spatial association) framework provides local versions for many spatial statistics (including point pattern and lattice data analysis) [anselinLocalIndicatorsSpatial1995].
-For example, Figure 2E shows LISA curves for the slice [getisSecondOrderNeighborhoodAnalysis1987, anselinLocalIndicatorsSpatial1995].
-That is, for each point, we calculate a local -function.
-These curves can be analysed with functional principal component analysis (fPCA), which allows to extract the main functional modes of variation [ramsayPrincipalComponentsAnalysis2005].
-The idea to perform fPCA on spatial statistics summaries has been used in ecological and spatial omics data [illianPrincipalComponentAnalysis2006, vuFunSpaceFunctionalSpatial2023, vuSPFSpatialFunctional2022].
-Figure 2F shows the scores of the first two functional PCs.
-There are two main clusters, representing either the physically clustered oligodendrocytes or those in low-density regions.
-Challenges and limitations
-Analysing biological samples formally as point patterns comes with several challenges and limitations.
-The first and most important is viewing the biological sample as the realisation of a point process.
-It is generally not easy to define whether the entire pattern of several cell types should be represented as a single point process (multitype analysis).
-Analysing the cell types individually and considering them as separate point processes might underestimate dependencies among the cells (multivariable analysis) [baddeleySpatialPointPatterns2015, pp.
-565 ff.].
-If, as in our example in Figure 2, the sections are far apart from each other, it can be justified to analyse the processes in isolation as a multivariable analysis.
-However, if we consider cells in two adjacent slices, these have probably been formed by one biological process and should not be analysed in isolation.
-Obtaining a representative sample of a point pattern is also a challenge.
-In spatial omics, we are often provided with FOVs that an experimentalist has selected from a much larger region or that stem from the technology itself.
-Such a sample might have different characteristics compared to the point pattern of the larger region, as FOVs are sometimes selected based on morphological properties (e.g., H&E staining).
-In cases where we have a small world scenario, we would want to limit our observation window to the full region where the points can occur.
-Lattice Data Analysis
-Definitions
-In contrast to point pattern-based methods, analysts could take the observation-based view and regard the location of cells or spots as a fixed (regular or irregular) lattice and consider the corresponding marker expression at each cell or spot.
-Several HTS-based spatial transcriptomics technologies produce data on a regular lattice (i.e., evenly spaced spots or beads of uniform size and shape), whereas imaging-based technologies yield irregular lattice structures (i.e., segmentation resulting in variable cell sizes and shapes, and non-uniform spacing) [zuurAnalysingEcologicalData2007, pp.
+Homogeneity (see Box Importance of scale and uniformity in point pattern analysis) assumes that the number of points in a given region is proportional to its area, i.e., homogeneity refers to a uniform intensity of points across the window of measurement.
+Otherwise, the point process is inhomogeneous [baddeleySpatialPointPatterns2015, p.132-133].
+This difference has important implications for the analysis of a point process.
+For example, when adjusting for varying local intensities, a process could be interpreted as inhomogeneous rather than genuinely clustered.
+This is called the confounding between intensity and interaction (see Box Importance of scale and uniformity in point pattern analysis) [baddeleySpatialPointPatterns2015, pp.151-152].
+Figure 4A shows a single FOV with islets from the human pancreas acquired using IMC [damondMapHumanType2019].
+An analysis considering the entire FOV will tell us about the processes that gave rise to the spatial distribution of the endocrine pancreas (islets).
+As we see in Figure 4B, a Ripley’s -function under a homogeneity assumption indicates that in this observation window, endocrine cells are clustered within pancreatic islets (the line lies above the theoretical CSR line).
+Figure 4B corrects for inhomogeneity of islets cells in the total fov and indicates CSR at .
+Note that the calculation of the inhomogeneous version varies with the estimation of the intensity function and the border correction (c.f., Supplementary Figure S5), especially when the data is not correlation-stationary [baddeleySpatialPointPatterns2015, pp.
+321ff.].
+However, there might also be a relevant biological scale where we subset our observation window to the islets themselves (Figure 4D).
+In this case, Ripley’s -function indicates a homogeneous distribution of the islet cells within the observation window (Figure 4E, the line follows the theoretical CSR line).
+Overall, the correct scale needs to be chosen in accordance with the research question, since the interpretation depends on it.
+Regular and irregular lattices
+Several HTS-based spatial transcriptomics technologies produce data on a regular lattice that gives evenly-spaced spots or beads of uniform size and shape, whereas imaging-based technologies yield irregular lattice structures with variable cell sizes and shapes, and non-uniform spacing [zuurAnalysingEcologicalData2007, pp.
 321ff.] [anselin2024introduction].
-Assumptions
 Once we have recorded the outline and arrangement of the spatial units in the lattice, we can specify the strength of spatial relationships between each unit.
 Each pair of units is assigned a weight: the stronger the connection between units, the higher the weight.
-The collection of all weights between locations and , , forms the weight matrix (see box Options for analysis).
-The construction of the weight matrix is critical for all downstream analyses as it encodes the spatial relationship between the units in the lattice [zuurAnalysingEcologicalData2007, pp.
-321ff.] and there exist various strategies: contiguity-based (i.e., in direct contact), graph-based, distance-based, or higher-order neighbours [getisSpatialWeightsMatrices2009, zuurAnalysingEcologicalData2007, pebesmaSpatialDataScience2023].
-Options for analysis
+The collection of all weights between locations forms the weight matrix (see Box Definition of the neighbourhood is critical in lattice data analysis).
+Spatial autocorrelation is a measure of similarity
 A common exploratory analysis for lattice data is based on the concept of spatial autocorrelation.
@@
 209ff.] [getisAnalysisSpatialAssociation1992].
-Over the years, various association measures have emerged, each specific to an aspect of the data.
-While many allow for univariate comparisons, in which one variable (e.g., expression of a gene) is compared over multiple locations, others allow for the comparison of two (bivariate) or more (multivariate) features at once [pebesmaSpatialDataScience2023, mosesVoyagerExploratorySinglecell2023].
-For most metrics, the features are assumed to be continuous measurements.
-However, metrics allow for the measurement of spatial autocorrelation of categorical variables (e.g., join count statistics).
-In addition to the number of features and the type of measurement that can be compared in the association score, there is the to consider: global spatial autocorrelation metrics estimate the average level of spatial autocorrelation across all locations whereas local measures give a statistic at each location.
-The global statistic can be seen as the weighted sum of its respective local statistics [anselinLocalIndicatorsSpatial1995].
-Example
-Moran’s is a prominent example of a spatial autocorrelation metric.
-It can be interpreted as the mean Pearson correlation between values at locations and the mean value of the corresponding neighbours [moranNotesContinuousStochastic1950].
-A positive value of Moran’s indicates that locations with similar values are clustered while a negative value indicates the clustering of dissimilar values, giving a measure of spatial heterogeneity.
-The global value of Moran’s is bounded by and with an expected value under spatial randomness near for large , ().
-Figure 3A shows evidence for spatial autocorrelation of Nrgn expression in mouse brain tissue; the (global) Moran’s value is (p-value essentially zero).
-Global measures can be used to identify spatially variable genes when compared to the global autocorrelation values of different genes [weberNnSVGScalableIdentification2023, chenEvaluatingSpatiallyVariable2024].
-Furthermore, there is a local version of Moran’s (interpretation is analogous to the global version).
-Figure 3B shows the estimate of local Moran’s at each measured spot (n.b., the local version is not bounded by and ).
-Figure 3C shows the corresponding adjusted p-values (testing for non-zero spatial autocorrelation).
-When interpreting local autocorrelation measures, it is important to consider both the effect size estimates and the significance level.
+When interpreting (local) autocorrelation measures, it is important to consider both the effect size estimates and the significance level.
 Since the significance level is calculated for each spot separately, it is recommended to adjust for multiple testing (e.g., Benjamini and Hochberg [benjamini_hochberg1995fdr]).
 Overall, local Moran’s statistics reveal locations that have similar values to their neighbours.
-However, as can be seen in Figure 3A-C regions with the highest and most significant local Moran’s value lie in the part of the tissue where expression of gene Nrgn is very low amongst neighbours.
-Notably, the local Moran’s measure is both dependent on the value of the log-transformed counts and the similarity among neighbours (c.f., Supplementary Figure S2A-B).
-For the analysis of categorical data, there are metrics such as the join count statistics [daleSpatialAnalysisGuide2014, pp.
-141ff.].
-In essence, the join count statistic calculates the frequency of categories among neighbours and compares this value with a theoretical distribution or permutations of the labels to get a significance score.
-This can be used to investigate the interactions between cell types in a lattice since the distance can be specified when constructing the neighbourhood graph.
-Challenges and limitations
-The choice of the weight matrix poses a challenge in the analysis of lattice data, as it might influence downstream analyses and conclusions.
-A common choice is to give a non-zero weight only to direct neighbours of a spatial unit (i.e., contiguity-based).
-Neighbours based on graphs or on distance [pebesmaSpatialDataScience2023, pp.
+However, in mouse brain, regions with the highest and most significant local Moran’s value lie where expression of gene Nrgn is very low among neighbours (Figure 5A-B).
+Notably, the local Moran’s measure is both dependent on the value of the log-transformed counts and the similarity among neighbours (c.f., Supplementary Figure S4A-B).Metrics such as the Moran’s scatter plot can be useful to classify the type of interaction [anselin2019moran].
+The Moran’s scatter plot compares the values of each spot with the average of it’s neighbours as defined in the weight matrix.
+This information can be used to classify each point into respective categories (c.f., Supplementary Figure S4E)).
+For example, in Figure 5 C the clusters denoted “high-high” refer to regions where high expression values are surrounded by high values, while “low-low” clusters indicate that low values are surrounded by low values (for which the local Moran’s value can be high as well).
+“High-low” and “low-high” values denote heterogeneous regions.
+Definition of the neighbourhood is critical in lattice data analysis
+The construction of the weight matrix is critical for all downstream analyses as it encodes the spatial relationship between the units in the lattice [zuurAnalysingEcologicalData2007, pp.
+321ff.].
+Several construction strategies exist: contiguity-based (i.e., in direct contact), graph-based, distance-based, or higher-order neighbours [getisSpatialWeightsMatrices2009, zuurAnalysingEcologicalData2007, pebesmaSpatialDataScience2023].
+Neighbours based on graph adjacency or on distance [pebesmaSpatialDataScience2023, pp.
 191 ff.] are also used; the open question is which is best suited for spatial omics data.
 If one uses contiguity-based neighbours in imaging-based spatial transcriptomics, results ultimately depend heavily on the accuracy of the cell segmentation, which is known to be challenging [stringerCellposeGeneralistAlgorithm2021, greenwaldWholecellSegmentationTissue2022, petukhovCellSegmentationImagingbased2022, architSegmentAnythingMicroscopy2023, wangCellSegmentationImage2019].
-If neighbours beyond the adjacent ones are used, the scale (e.g., at which cells interact still) needs to be specified.
+If neighbours beyond the adjacent ones are used, the scale (e.g., at which cells interact) needs to be specified.
 Figure 5C-D show the difference between the local Moran’s calculation when based on contiguity-based neighbours in D, the 10 nearest neighbours in E or neighbours within a 1000 pixel distance ( ) in F.
@@
 Smoothing of the local Moran’s values occurs when more neighbours are considered (Supplementary Figure S4C).
-Supplementary Figure S2A and B further show that for some cells, no contiguous neighbours were found, which results in a zero estimate of the local Moran’s .
-Since cells do not function in isolation but form complex anatomical structures including extracellular components, an open question is whether analyses improve when weight matrix construction takes anatomical structures or regions into account.
-Overall, it remains to be investigated how much the construction of the weight matrix influences downstream analyses in spatial omics data.
-Short guide to spatial analysis of biological data
-The large variety of options for analysis in spatial statistics can be challenging to navigate.
-A few fundamental steps should help selecting appropriate exploratory analysis.
+Furthermore, Supplementary Figure S4A and B show that for some cells, no contiguous neighbours were found, which results in a zero estimate of the local Moran’s .
+Cells do not function in isolation but form complex anatomical structures including extracellular components.
+Whether analyses improve when weight matrix construction takes such anatomical structures or regions into account is to be determined.
+Furthermore, it remains to be investigated how much the construction of the weight matrix influences downstream analyses in spatial omics data.
+pasta highlights various spatial statistics analyses with inline code in R and Python
+The collection of vignettes on our webpage (https://robinsonlabuzh.github.io/pasta) showcases the analysis of data from multiple spatial omics technologies, where concepts and assumptions are discussed in detail, with inline R and Python code.
+The design of the webpage follows the outline of Figure 1.
+First, we consider the technology and categorize into a HTS-based or imaging-based approach.
+Second, we distinguish between univariate (one type of mark) and multivariate (more than one type of mark) methods and outline different methods in specific vignettes with an in-depth discussion of the methods, e.g.
+providing mathematical definitions and formulae.
+In addition, we provide overview vignettes summarizing the two approaches.
+To ensure usability, we showcase exploratory spatial statistics methods integrated into existing bioinformatics ecosystems.
+In Python, we use anndata objects from scanpy, allowing direct usage with SpatialData as well [virshup2024anndata, wolf2018scanpy, marconato2025spatialdata].
+In R, we integrate with SpatialExperiment and SpatialFeatureExperiment objects from Bioconductor [righelliSpatialExperimentInfrastructureSpatiallyresolved2022, mosesVoyagerExploratorySinglecell2023, gentleman2004bioconductor].
+The packages Voyager and spatialFDA provide wrappers to compute both lattice and point pattern metrics on Bioconductor objects [mosesVoyagerExploratorySinglecell2023, emonsSpatialFDAToolSpatial2025].
+Furthermore, our vignette offers examples how to convert between anndata objects in Python and SpatialExperiment, Seurat objects objects in R [righelliSpatialExperimentInfrastructureSpatiallyresolved2022, haoDictionaryLearningIntegrative2023].
+Using these conversions, we are enabling analysis for users of both R and Python.
+Overall, our resource Pattern Analysis for SpaTial omics datA (pasta) highlights the usefulness and transferability of existing exploratory spatial statistics approaches in the context of spatial tissue profiling.
+Spatial statistics is used in existing spatial omics data analysis tools
+There are already several tools for omics analysis that employ spatial statistics.
+Among others, Voyager is a framework and a collection of use cases for lattice data written in both R and Python [mosesVoyagerExploratorySinglecell2023]; it provides data structures and functionalities to compute spatial statistics in molecular biology.
+Voyager comes with comprehensive vignettes for several spatial profiling technologies.
+The methodological focus of their vignettes is on lattice data analysis and the spdep geospatial package [bivandPackagesAnalyzingSpatial2022].
+Since Voyager offers efficient implementations for lattice data analysis in R, we build on their framework in our vignettes for the lattice data component.
+Other methods written in R include: SPIAT, which contains various tools for spatial analysis, such as: neighbourhood analysis, local metrics and heterogeneity scores etc.
+[fengSpatialAnalysisSPIAT2023]; scFeatures, a toolbox comprising lattice data analysis functionality such as Moran’s and point pattern metrics as the -function [caoScFeaturesMultiviewRepresentations2022]; spicyR, which uses a compressed version of the global -function for cross-sample comparison [caneteSpicyRSpatialAnalysis2022]; lisaClust, a spatial domain detection method via LISA (local indicators of spatial association) curves [patrickSpatialAnalysisHighly2023]; mxfda, that uses point pattern summary functions for survival analysis [wrobelMxfdaComprehensiveToolkit2024]; spatialDM uses global and local bivariate Moran’s to score co-localization of ligand-receptor pairs [liSpatialDMRapidIdentification2023]; and MERINGUE, which uses nearest-neighbour autocorrelation for spatially variable gene selection [millerCharacterizingSpatialGene2021].
+Methods in Python include squidpy, a package with various spatial statistics tools for both point pattern and lattice data [pallaSquidpyScalableFramework2022]; PySAL a general Python library for spatial analysis [reyPySALPythonLibrary2010]; Bento a toolkit for analysis of subcellular features using different spatial statistics [mahBentoToolkitSubcellular2024]; SpottedPy a package designed to identify tumor hotspots using spatial statistics [withnellSpottedPyQuantifiesRelationships2024]; MuSpAn a toolbox for multiscale spatial analysis using spatial statistics [bullMuSpAnToolboxMultiscale2024] and spatiomic a Python library for the analysis of pixel based spatial data using spatial correlation measures [kuehlSpatiomicScalableSpatial2025].
+For a detailed view on general methods development in the field of spatial transcriptomics, we refer readers to the online Supplementary table of the museum of spatial transcriptomics paper [mosesMuseumSpatialTranscriptomics2022].
+Short guide to spatial analysis of spatial omics data
+Step 1 – The data representation
 First, analysts should consider their technology and what options they have (i.e., imaging-based or HTS-based approaches).
@@
 In point pattern analysis, the process leading to the locations of individual points is the primary interest.
-It allows us to make statements such as whether a point process exhibits clustering or if two cell types tend to interact more than randomly expected.
-In lattice data analysis, however, we are not directly interested in the arrangement of points, but rather locations are important to define the spatial relationship between locations (i.e., via the weight matrix).
-We study the interaction of features, such as gene expression, given their spatial location.
-For example, if ligand-receptor dynamics are of interest, the weight matrix can be specified in a way that captures spatial associations of gene expression between direct neighbours or at a specified scale.
-Second, the observation scale has to be chosen.
-Sometimes, this is a technological scale such as multiple FOVs in imaging.
-Figure 4A shows a single FOV with islets from the human pancreas acquired using IMC [damondMapHumanType2019].
-An analysis considering the entire FOV will tell us about the processes that gave rise to the spatial distribution of the endocrine pancreas (islets).
-As we see in Figure 4B, a Ripley’s -function under a homogeneity assumption indicates that in this observation window, endocrine cells are clustered within pancreatic islets.
-However, there might also be a relevant biological scale where we subset our observation window to the islets themselves (Figure 4D).
-In this case, Ripley’s -function indicates a homogeneous distribution of the islets cells within the observation window (Figure 4E).
-Overall, the correct scale needs to be chosen in accordance with the research question, since the interpretation depends on it.
-In point pattern analysis, the analyst needs to decide whether (and how) to control for inhomogeneous intensities, leading back to potential confounding of intensity and interaction [baddeleySpatialPointPatterns2015, pp.
-151 ff.].
-Analysts can look at smoothed density estimates to inspect potential inhomogeneity prior to analysing the pattern more formally.
-Furthermore, it depends again on the research question and biological domain knowledge whether the assumption of homogeneity is appropriate.
-While tissues such as the adipose tissue form regular structures that could be assumed to be homogeneous, other tissues such as stratified epithelial structures appear to be more inhomogeneous.
-As shown in Figure 4, the homogeneous and inhomogeneous variants of the spatial summaries can lead to different interpretations.
+It allows us to make statements such as whether a point process exhibits clustering or if two cell types tend to interact more than expected under CSR.
+In lattice data analysis, however, we are not directly interested in the arrangement of points, but rather locations are used to define the structure of the neighbourhood (i.e., via the weight matrix).
+We study the interaction of features, such as gene expression levels, given their spatial location.
+Step 2 – Choice of feature and nature of the spatial interaction
+Second, the feature has to be considered.
+Most often this will be either gene or protein expression (continuous feature) or cell type or other labels (categorical feature).
+As seen in Figure 2, lattice data analysis offers many more methods for continuous features, whereas the focus of point pattern analysis is categorical features.
+Next, the scientific question guides the choice of the analysis.
+For example, if we are interested in the expression of one gene in space, this is a univariate continuous comparison.
+Another example would be the comparison of the co-localization of two cell types, which is a bivariate categorical comparison.
 A fundamental choice in lattice data analysis is the construction of the weight matrix.
-As seen in Figure 3A-B, there are differences in the interpretation of results between different weight matrix choices.
+As seen in Figure 5D-F, there are differences in the interpretation of results between different weight matrix choices.
 The appropriate weight matrix design depends on the research question.
@@
 If the feature of interest is a molecule that diffuses to nearby cells (e.g., paracrine signalling), a nearest neighbour or distance-based approach should be considered for constructing the neighbourhood matrix.
+Figure 1 helps with the navigation of both data modalities and analysis options.
+Step 3 – Observation scale
+Third, the observation scale has to be chosen.
+Sometimes, a scale given by a technology such as a FOV might be appropriate.
+In other cases, the observation scale should be given by the biological question.
+Recall the example of Figure 4 where the scale effectively determines whether the point pattern of islet cells is considered to be homogeneous or clustered.
+In lattice data analysis the scale is directly connected to the weight matrix.
+For example, if ligand-receptor co-localizations are of interest, the weight matrix can be specified in a way that captures spatial associations of gene expression between direct neighbours or at a specified scale.
+Generally, the concept of spatial autocorrelation is based on Tobler’s famous first law of geography, which states that “everything is related to everything else, but near things are more related than distant things” [toblerComputerMovieSimulating1970].
+This statement is not generally applicable in biology where we face highly compartmentalized anatomical structures.
+In analysis, it might therefore be important to account for this compartmentalization and define the neighbourhood accordingly or perform analysis that is aware of different compartments.
+Step 4 – Technical considerations
+In point pattern analysis, the analyst needs to decide whether (and how) to control for inhomogeneous intensities, leading back to potential confounding of intensity and interaction [baddeleySpatialPointPatterns2015, pp.
+151 ff.].
+Analysts can look at smoothed density estimates to inspect potential inhomogeneity prior to analyzing the pattern more formally.
+Furthermore, it depends again on the research question and biological domain knowledge whether the assumption of homogeneity is appropriate.
+While some tissues form regular structures that could be assumed to be homogeneous, other tissues appear to be more inhomogeneous.
+As shown in Figure 4, the homogeneous and inhomogeneous variants of the spatial summaries can lead to different interpretations.
 Overall, we advise analysts to consider both scale and homogeneity carefully, keeping the biological context in mind and to document the decisions made and their rationale.
 Discussion
-In this work, we presented multiple options for the exploratory data analysis of spatial omics data.
+In this work, we presented multiple options for the exploratory analysis of spatial omics data.
 We introduced the foundations of both point pattern and lattice data analysis and showed the usability of both frameworks in accompanying vignettes.
-Fundamentally, point pattern and lattice data analysis are two distinct paths to analysing spatial omics data.
+Fundamentally, point pattern and lattice data analysis are two distinct paths to analyzing spatial omics data.
 Analysts should be aware of the assumptions and interpretations that can be drawn from each approach.
@@
 In contrast, point pattern analysis concentrates on the underlying mechanisms that generate stochastic spatial patterns (i.e., clustered, regular or spaced arrangements).
+Moreover, the type of mark associated with spatial locations is another decisive factor in analysis.
+Lattice data analysis is most suited when the mark of interest is continuous, while point pattern analysis is the case for categorical marks.
+However, small nuances exist with lattice data analysis methods for categorical data and point pattern methods for continuous data.
 Therefore, the choice of the analysis framework depends not only on the data modalities, but also on the research question [anselin2024introduction, cressieStatisticsSpatialData1993, baddeleySpatialPointPatterns2015, pebesmaSpatialDataScience2023, daleSpatialAnalysisGuide2014].
 Meanwhile, there are dualities between the data modalities from spatially-resolved measurements (c.f., Figure 1).
-Using cell segmentation, imaging data can be approximated as a points via their centroids.
-Point locations are stochastic and can therefore be regarded as a realisation of a point process and can be studied using point pattern analysis.
-On the other hand, we can also view segmentations as an irregular lattice (and aggregate expression per cell).
-Similarly, HTS-based data with subcellular resolution (e.g., Visium HD [oliveiraCharacterizationImmuneCell2024]) can likewise be interpreted as a regular lattice of (sub)-cellular locations or irregular lattice from segmentations [polanskiBin2cellReconstructsCells2024] or the reconstructed cells can then be approximated by their centroids as points and further analysed using point pattern analysis.
+Using cell segmentation, imaging data can be approximated as points via their centroids.
+Point locations are stochastic and can therefore be regarded as a realization of a point process and can be studied using point pattern analysis.
+On the other hand, we can also view segmentation as an irregular lattice (and aggregate expression per cell).
+Similarly, HTS-based data with subcellular resolution (e.g., Visium HD [oliveiraCharacterizationImmuneCell2024]) can likewise be interpreted as a regular lattice of (sub)-cellular locations or irregular lattice from segmentation [polanskiBin2cellReconstructsCells2024] or the reconstructed cells can then be approximated by their centroids as points and further analyzed using point pattern analysis.
 We have focused here on spatial statistics and highlighted two main streams of analysis.
-Another popular way to represent spatial arrangements is spatial graphs [aliGraphCompassSpatialMetrics2024, windhagerEndendWorkflowMultiplexed2023].
+Another popular way to represent spatial arrangements is as spatial graphs [aliGraphCompassSpatialMetrics2024, windhagerEndendWorkflowMultiplexed2023].
 The methods described in the section on lattice data are closely connected to spatial graphs, especially in the construction of a weight matrix.
-Point pattern analysis concepts are also related to spatial graph via edge rules.
-For example, Ripley’s can be interpreted as a graph problem with a distance threshold edge rule, as explained in [daleSpatialAnalysisGuide2014, p.
+Point pattern analysis concepts are also related to spatial graphs via edge rules.
+For example, Ripley’s can be interpreted as a graph problem with a distance threshold edge rule [daleSpatialAnalysisGuide2014, p.
 136].
+Finding the correct scale of analysis
 One common challenge in all spatial analyses is the question of the correct scale.
@@
 […]
+Limitations of exploratory spatial statistics
 There are various other limitations when studying spatially-resolved omics data.
 On the technical side, imperfect sections can lead to artifacts or non-overlapping FOVs, which can introduce missing data.
-Furthermore, cells are never perfectly arranged within a histological section, thus cells overlapping in the z-axis are to be expected.
+Furthermore, cells are never a perfectly monolayer within a histological section, thus cells overlapping in the -axis are to be expected.
 In terms of analysis, two-dimensional views neglect processes that occur in three dimensions.
@@
 Both lattice data and point pattern analysis extend conceptually to 3D, however with increased computational complexity.
-Moreover, in point pattern analysis, we often implicitly assume that patterns are invariant to rotations (see box Assumptions).
+Moreover, in point pattern analysis, we often implicitly assume that patterns are invariant to rotations (see Box Importance of scale and uniformity in point pattern analysis).
 This assumption is not anatomically accurate for organs and tissues with layered structures, such as the brain.
-There are other limitations arising from the concept of spatial metrics.
-Lattice data analysis methods were originally designed for rather low dimensional datasets, while spatial omics often deals with thousands of features (e.g., full transcriptome).
-Similarly, point pattern methods usually only allow for pairwise comparison of features, which can lead to a high number of cross comparisons e.g., when there are a dozen cell types.
-Moreover, with newer technologies, the size of the biological sample and resolution of measurements are constantly expanding, which can lead to computational performance issues (e.g., to store cell segmentations in memory to construct weight matrices).
+Moreover, there are other limitations arising from the concept of spatial metrics.
+Lattice data analysis methods were originally designed for rather low dimensional datasets, while spatial omics often deals with thousands of features (e.g., untargeted transcriptome).
+In addition, the interpretation of a spatial correlation metric depends on the measure of similarity that is compared in a local neighbourhood.
+In a multivariate setting, a straightforward interpretation of this similarity measure is often not possible.
+Similarly, point pattern methods usually only allow for pairwise comparison of features, which can lead to a high number of cross comparisons.
+Moreover, with newer technologies, the size of the biological sample and resolution of measurements are constantly expanding, which can be computationally expensive.
 Since point pattern based methods only store coordinates, they usually scale better in terms of memory requirements.
-In our accompanying vignette, we discuss concepts and assumptions in more detail using biological examples with R and Python.
+In our accompanying vignette, we discuss concepts and assumptions in more detail using biological examples.
 For point pattern analysis, we mainly use R due to the lack of comprehensive Python libraries available for point pattern analysis.
@@
 In addition, we provide vignettes for point pattern analysis to complement existing discussions [zormpasMappingTranscriptomeRealizing2023, mosesVoyagerExploratorySinglecell2023].
-Outlook
-The spatial metrics as described in point pattern and lattice data analysis offer interesting ways to quantify spatial patterns in molecular biology.
+Future opportunities
+The spatial metrics as described in point pattern and lattice data analysis offer means to quantify spatial patterns in molecular biology.
 There are still open research questions including the correct application (e.g., scale and homogeneity) of methods to a given biological context, the high dimensional nature of spatial omics data and the effect of preprocessing steps that come upstream of the application of spatial statistics.
 Furthermore, comparisons of spatial metrics across multiple samples and conditions and parametric spatial models offer interesting options for future work.
+Methods
+All analysis were conducted using R version 4.5.0 and Bioconductor version 3.21 [gentleman2004bioconductor]. Results were visualized using ggplot2 [wickhamGgplot2ElegantGraphics2016].
+[…]
+The mouse brain datasets shown in Figure 1 (Technology) were taken from Shi et al.
+(STARMAP Plus) [shiSpatialAtlasMouse2023] [shiSpatialAtlasMouse2023] and 10x (Visium mouse) [10xVisium_Mouse].
+Clustering was performed using Banksy [singhalBANKSYUnifiesCell2024].
+Lattice data analysis was performed using the package Voyager [mosesVoyagerExploratorySinglecell2023] and spdep [pebesmaAppliedSpatialDataAnalysis2013], point pattern analysis was conducted using the packages spatialFDA [emonsSpatialFDAToolSpatial2025] and spatstat [baddeleySpatstatPackageAnalyzing2005].
+Biologically relevant observation windows were determined using the package sosta [gunzSostaPackageAnalysis2025].
 Code and data availability
-Code and instructions to retrieve data to reproduce figures and results can be found under https://github.com/robinsonlabuzh/pasta-manuscript, code licensed under the GPLv3 terms.
+Code and instructions to retrieve data to reproduce figures and other results can be found under https://github.com/robinsonlabuzh/pasta-manuscript, code licensed under the GPLv3 terms.
 The vignette and source code can be accessed via https://robinsonlabuzh.github.io/pasta licensed under CC-BY-4.0 terms for the text and GPLv3 license for any code.
+Versioned source code and rendered vignettes available under https://doi.org/10.5281/zenodo.15488040.
 Acknowledgments
 We thank all members of the Robinsonlab for constructive feedback on the manuscript and vignette.
-In particular, we thank Peiying Cai, Alice Driessen, Reto Gerber, Pierre-Luc Germain, Maruša Kodermann, Vladyslav Korobeynyk, Siyuan Luo, Giulia Moro, Emanuel Sonder, Jiayi Wang and David Wissel for their careful reading and input to the vignette.
-The Python examples in the vignette were developed as part of the SpaceHack 2024 event.
+In particular, we thank Peiying Cai, Alice Driessen, Reto Gerber, Pierre-Luc Germain, Maruša Kodermann, Vladyslav Korobeynyk, Siyuan Luo, Giulia Moro, Emanuel Sonder, Jiayi Wang and David Wissel for their careful reading and input to the vignette.The Python examples in the vignette were developed as part of the SpaceHack 2024 event.
 We thank the SpaceHack organisers and participants.
+Furthermore, we would like to thank Daria Lazic for helpful comments on the manuscript.
 Author contributions
-ME: Conceptualisation, Methodology, Software (Vignette), Analysis, Visualisation, Writing - Original Draft; SG: Conceptualisation, Methodology, Software (Vignette), Analysis, Visualisation, Writing - Original Draft; HLC: Conceptualisation, Vignette, Writing - Review & Editing; IM: Methodology, Writing - Review & Editing; MK, Software (Vignette), Writing - Review & Editing; RF: Methodology, Writing - Review & Editing; MDR: Conceptualisation, Methodology, Writing - Original Draft, Supervision, Funding acquisition.
+ME: Conceptualisation, Methodology, Software (Vignette), Analysis, Visualisation, Writing - Original Draft; SG: Conceptualisation, Methodology, Software (Vignette), Analysis, Visualisation, Writing - Original Draft; HLC: Conceptualisation, Vignette, Writing - Review & Editing; IM: Methodology, Software (Vignette), Writing - Review & Editing; MK, Software (Vignette), Writing - Review & Editing; RF: Methodology, Writing - Review & Editing; MDR: Conceptualisation, Methodology, Writing - Original Draft, Supervision, Funding acquisition.
 Funding
@@
 We declare no conflict of interest.
-Supplementary Figures
+Supplementary data
+Technology | Main feature of interest | Lattice data | Point pattern | |||
+Visium [stahlVisualizationAnalysisGene2016] | transcriptomics | spots (regular) | not applicable | |||
+Slide-seq V1& V2 [rodriquesSlideseqScalableTechnology2019, stickelsHighlySensitiveSpatial2021] | transcriptomics |
+segmented cells | |||
+Visium HD [oliveiraCharacterizationImmuneCell2024] | transcriptomics |
+segmented cells | |||
+Stereo-seq [chenSpatiallyResolvedHighly2015] | transcriptomics |
+segmented cells | |||
+Xenium [janesickHighResolutionMapping2023] | transcriptomics | segmented cells (irregular) |
+CosMx [chenSpatiallyResolvedHighly2015] | transcriptomics | segmented cells (irregular) |
+MERFISH [chenSpatiallyResolvedHighly2015] | transcriptomics | segmented cells (irregular) |
+IMC [giesenHighlyMultiplexedImaging2014] | proteomics |
+segmented cells | |||
+MIBI-TOF [kerenMIBITOFMultiplexedImaging2019] | proteomics |
+segmented cells | |||
+CODEX [blackCODEXMultiplexedTissue2021] | proteomics |
+4i [gutMultiplexedProteinMaps2018] | proteomics |
+![[Uncaptioned image]](x9.png)
+(Figure legend on next page.)